OpenStudy (anonymous):

Can this be integrated: ln t /[t * e^(4t) ]

OpenStudy (beginnersmind):

This isn't really an answer but my idea would be to write it as (ln t/t)*e(-4t). The first part is in the form f'(x)*f(x) so it can be integrated. So integration by parts should work.

OpenStudy (anonymous):

But even if I can integrate the first part, how does that help me? The function needs to be integrated as a whole. Maybe I'm misunderstanding what you tried to say.

OpenStudy (anonymous):

If you write it like this, you could do integration by parts: \[\int\limits \ln(t) * (1/te ^{4t}) dt\]

OpenStudy (anonymous):

The integration by parts formula is \[\int\limits u dv = uv-\int\limits v du\] so choose either ln(t) or (1/te^4t) to be u and the other to be dv, and use the formula. I don't know what kind of answer that gives as I haven't done it myself but that seems to be a logical approach.

OpenStudy (anonymous):

Or if you go by what beginnersmind is saying, you could write it like that too and do integration of parts but for ln(t)/t you'd have to use the product rule inside your integration by parts.

OpenStudy (anonymous):

Thanks, yeah that makes sense.

OpenStudy (beginnersmind):

Actually the trick is that the derivative of (lnx)^2 is 2lnx/x. So the right way to separate seems to be dv = lnx/x and u = e^(-4t)

OpenStudy (anonymous):

Thanks to both of you, I'm gonna try it out.

OpenStudy (beginnersmind):

Actually, just ignore everything I said :( I don't think it's helpful at all...

OpenStudy (anonymous):

I've tried it several ways now and all ways look complicated. It looks like you'd just have to keep doing integration by parts forever. There must be a trick somewhere to simplify it...

OpenStudy (anonymous):

yeah this seems like a very complicated integral...round and round